OpenStudy (anonymous):

Try for fun. Solve for x, Given that x^4-2x^3-5x^2+10x-3=0.

OpenStudy (anonymous):

It gonna have 4 solutions..

OpenStudy (anonymous):

ok, then,

OpenStudy (experimentx):

the easiest way is to factorize it .. into lower terms. best of luck @Eyad

OpenStudy (anonymous):

if you factorize it to the lowest power of x^2 you'll get x=3 or -7 or (2+\[\sqrt{24})/2\]or (2-\[\sqrt{24})/2\]

OpenStudy (maheshmeghwal9):

\[x^4-2x^3-5x^2+10x-3=0.\]

OpenStudy (maheshmeghwal9):

now looking good:)

OpenStudy (anonymous):

thank u @maheshmeghwal9

OpenStudy (maheshmeghwal9):

it's really a fun:)

OpenStudy (maheshmeghwal9):

yw:)

OpenStudy (anonymous):

grr, no rational roots.. Descartes' rule shows 3 positive roots and 1 negative root. From plotting a few points, there are roots -3<x<-2, 0<x<1, 1<x<2, and 2<x<3. Going to have to bring out the big guns to find those irrationals. ;-)